OpenStudy (anonymous):

Explain why the function g(x) = x|x| has and inflection point at (0, 0) even though g"(0) does not exist. I know that g' is 0, and there is no rate of change for it and that's why g" DNE. But i can't see why it has an inflection point.

OpenStudy (perl):

$$ \Large g(x) = x\cdot |x| $$

OpenStudy (anonymous):

Yes it is.

OpenStudy (perl):

note that $$ \\\Large |x|'= \frac {x}{|x|} $$

OpenStudy (anonymous):

Oh so looking at g'(x) we can see that there is a inflection point at x=0 if you set the numerator to zero?

OpenStudy (perl):

well you need to find the second derivative, and then find where g ' ' (x) = 0 or is undefined . we see that it is undefined at x = 0. now this is a potential inflection point. to show it is a true inflection point , show that the concavity changes. that meanst f ' ' (x) changes sign about x = 0

OpenStudy (perl):

you can use quotient rule to find g ' ' (x)

OpenStudy (perl):

$$ \Large { g(x)= x \cdot |x| \\ \text {product rule} \\ \Large g ' (x) = 1 \cdot |x| + x \cdot \frac {x}{|x|} = \frac{ |x|\cdot |x| }{|x|} + x \cdot \frac {x}{|x|}\\= \frac{x^2+x^2}{|x|}= \frac{2x^2}{|x|}\\ \\ \therefore \\ g '' (x) = (\frac{2x^2}{|x|}) ' = \frac{|x|\cdot 4x -2x^2\cdot \frac{x}{|x|} }{|x|^2} = \frac {2x}{|x|} } $$

OpenStudy (anonymous):

So i can prove that there is an infliction by actually finding it?

OpenStudy (perl):

yes, we can actually find the inflection point. the inflection point is where g ' ' (x) = 0 , or is undefined. (assuming this x is in the domain of g(x))

OpenStudy (perl):

and g(x) is defined for x = 0. g(0) = 0 * |0 | = 0

OpenStudy (perl):

now to prove this is a true inflection point, test g ' ' ( -1) and g ' ' ( 1) . if the sign of the second derivative changes, then x = 0 is a true inflection point

OpenStudy (anonymous):

I got an answer that shows it does change. By looking at this how can i tell if it is concave up or down?

OpenStudy (anonymous):

for -1 i got a negative 2nd derivative and for +1 i got a positive deriv, does that mean it is concave up?

OpenStudy (anonymous):

any time it doesn't exist then the inflection point is (0,0)